Victor spent 3/8 of his money in one of his shop, 1/4 in the second shop and 1/8 in the third shop. He then had #100 left. How much did he start with?
Answer by greenestamps(13200)



3/8 + 1/4 + 1/8 = 3/8 + 2/8 + 1/8 = 6/8 = 3/4

He has spent 3/4 of his money; he has 1/4 of it left. Since 1/4 of his money is #100, the amount he started with is #400.